OpenStudy (anonymous):

Determine the exact value: cos(sin^-1(1/2))

myininaya (myininaya):

i will draw you a picture

OpenStudy (anonymous):

Thanks that'd be great

myininaya (myininaya):

i always find it useful to draw a right triangle to do these problems

OpenStudy (anonymous):

draw a triangle. put one angle as theta say. then using "opposite over hypotenuse" put opposite as 1 and hypotenuse as 2. find that the third side is \[\sqrt{3}\] by pythagoras so \[\cos(\theta)=\frac{\sqrt{3}}{2}\]

myininaya (myininaya):

do you see how i found the missing side i used the pythagoren thm

OpenStudy (anonymous):

thats a beautiful triangle

myininaya (myininaya):

don't make fun of my triangle

OpenStudy (anonymous):

that is method one. method two is to say \[\sin^{-1}(\frac{1}{2})=\frac{\pi}{6}\] and \[\cos(\frac{\pi}{6})=\frac{\sqrt{3}}{2}\]

myininaya (myininaya):

lol

OpenStudy (anonymous):

a picture is worth a thousand ships.

OpenStudy (anonymous):

That really helps actually. You wouldn't happen to have the full chart handy would you?? I have like 6 more problems with sin/cos/tan

myininaya (myininaya):

the unit circle?

myininaya (myininaya):

you can find some pretty unit circles online

OpenStudy (anonymous):

yes that's what it's called! i couldn't remember. but i couldn't search it till i knew what it was, so thanks lol

myininaya (myininaya):

np :)

OpenStudy (anonymous):

here is a nice one

OpenStudy (anonymous):

got all the stuff. unit circle on last page

myininaya (myininaya):

i was kidding